OpenStudy (anonymous):

Which of the following is equivalent to 24 =16? log(base 4)16 = 2 log(base 2)4 = 16 log(base 4)2 = 16 log(base 2)16 = 4 Which of the following is equivalent to 24 =16? log(base 4)16 = 2 log(base 2)4 = 16 log(base 4)2 = 16 log(base 2)16 = 4 @Mathematics

OpenStudy (ash2326):

first one

OpenStudy (anonymous):

How come? because im stuck between the first and last

OpenStudy (ash2326):

oh yeah the last one is right

OpenStudy (anonymous):

Lol so which one is, and how come?

OpenStudy (ash2326):

log(base 2) 16=4 so 2^4=16 and that's our ques 2^4=16

OpenStudy (ash2326):

first one is 4^2=16

OpenStudy (anonymous):

first one right

OpenStudy (anonymous):

but both first and last are true

OpenStudy (anonymous):

I think it is the last one, since that is the logarithmic form of 2^4 = 16, no?

OpenStudy (anonymous):

but 4^2 is 16 too

OpenStudy (anonymous):

Oh wow, thats true, but according to the way its suppose to be changed: y = b^x <--> log(base b) y = x

OpenStudy (anonymous):

So it would be the last one with knowing that ^